# AstaGen: Empirical Evaluation of Automated SATD Taxonomy Generation with LLMs

### v1

@@ IV-E3 RQ3
For each dataset, we calculate the average execution time and cost of running AstaGen ten times. In AstaGen, the cost only arises from the use of the DeepSeek-V3 API. During each run, our implementation records the total number of input and output tokens, and estimates the API cost based on the official pricing table of DeepSeek-V3.555https://api-docs.deepseek.com/quick_start/pricing, as of May 20, 2025
@@ V Results
V-A RQ1: Category Name Alignment
Figure 4 shows the distribution of $top\_sim(H_{i})$ scores. The x-axis denotes the dataset, and colors indicate the taxonomy level. The mean $top\_sim$ scores for the QS, SC, and ML datasets are 0.36, 0.32, and 0.42, respectively.
AstaGen can generate domain-specific category names that are similar to human-defined ones. As highlighted in Figure 4, all points with scores above $\geq 0.8$ correspond to domain-specific categories (e.g., Layer Configuration in ML).
High-similarity categories are more likely to be domain-specific, whereas low-similarity categories tend to be generic. To examine the relationship between $top\_sim(H_{i})$ and the type of category, two authors independently labeled each human-defined category in the top and bottom 10% of $top\_sim(H_{i})$ as either domain-specific or generic, counting only those with agreement. Among the top 10% (77 categories), 46 ($\simeq$ 60%) were domain-specific, supporting the earlier observation. In the bottom 10%, 53 ($\simeq$ 69%) were labeled as generic. These results suggest that AstaGen more effective at generating consistent names for domain-specific categories, while generic categories are harder to match semantically.
Answer to RQ1: AstaGen tends to generate domain-specific category names that are semantically similar to human-defined names (e.g., Layer Configuration in ML). In contrast, low-similarity cases were predominantly generic (69% of 77 categories).
TABLE I: Comparison of AstaGen and its naive variant.
@@ A N A N A N A N
A: AstaGen, N: Naive. Bold indicates the better value.
V-B RQ2: Taxonomy Content Alignment
Table I presents the mean precision, recall, F1 scores, and differences in the number of categories for each dataset and taxonomy level. Although these metrics capture two aspects (i.e., comment assignment accuracy and category granularity), there are no guidelines for interpreting their absolute values. Therefore, we introduce a baseline, the naive variant, which generates the taxonomy using an LLM without Phases 1 and 2 of AstaGen. This baseline allows us to evaluate the extent to which AstaGen improves alignment with the human-defined taxonomy compared to a naive use of LLMs.
AstaGen achieves higher F1 scores than the naive variant in five out of six cases, indicating a better balance between precision and recall. This result suggests that core components of AstaGen, which the naive variant lacks, are effective in improving the alignment of comment assignments between human-defined and AstaGen-generated categories.
AstaGen also better matches the granularity of human-defined taxonomies in four out of six cases. In four out of six cases, AstaGen exhibits a smaller absolute difference than the naive variant, suggesting that its taxonomy size more closely matches human taxonomies. In contrast, the naive variant often produces fewer categories, which likely contributes to its higher recall. These findings indicate that AstaGen outperforms the naive variant by producing taxonomies that are not only more accurate in assignment but also more comparable in structural granularity. Hence, AstaGen generates similar categories to human-defined ones compared to the naive variant.
Answer to RQ2: AstaGen achieves higher F1 scores than the naive variant in five out of six cases. It also generates a number of categories more similar to human-defined taxonomies in four out of six cases. Hence, AstaGen generates categories that are more similar to human-defined ones compared to the naive variant. However, the absolute values of precision and recall remain between 0.152 and 0.638, indicating that there is still room for improvement in comment-level alignment.
TABLE II: Average execution time and cost of AstaGen for each dataset.
@@ V-C RQ3: Efficiency (Time and Cost)
Table II shows the average execution time and cost of AstaGen for each dataset, averaged over ten runs.
Compared to prior manual efforts, AstaGen achieves significantly reduces both time and cost. For example, Ebrahimi et al. [6] reported spending 57 person-hours just to construct the initial taxonomy, while Xiao et al. [8] required 120 person-hours to manually inspect 500 SATD comments. In contrast, AstaGen categorizes hundreds of comments in under two hours at a cost of less than one USD. These substantial efficiency gains highlight the practicality of using LLMs for large scale SATD taxonomy generation.
Answer to RQ3: AstaGen is efficient in terms of both time and cost. Even for the largest dataset we used (i.e., the ML dataset with 448 comments), it took only 109.50 minutes and cost just 0.813 USD on average.
VI Use Case: Semi-Automated Taxonomy Generation
Although AstaGen may not fully automate SATD taxonomy generation—as RQ1 and RQ2 showed that its alignment with human-defined taxonomies was not always high—it remains a practical assistant for reducing human effort. We propose two use cases for AstaGen: (1) as a virtual annotator that can replace some human annotators in collaborative taxonomy construction, and (2) as an initial step that provides category candidates to support annotators. These use cases support semi-automated taxonomy generation. By incorporating these human-in-the-loop workflows, it becomes possible to balance the efficiency of automation (as demonstrated in RQ3) with the quality of the resulting taxonomy through human review.
Notably, AstaGen has demonstrated the ability to generate domain-specific categories. This makes it especially valuable when researchers must construct a taxonomy in an unfamiliar domain. By incorporating AstaGen as one of the contributors, researchers can accelerate the early stages of taxonomy construction. In addition to improving AstaGen, evaluating these use cases remains future work.
@@ VII Threats to Validity
Construct Validity. In RQ2, we evaluate the alignment of AstaGen-generated taxonomies with human-defined ones using best-match precision and recall. This evaluation is conducted in a relative manner using the naive variant, and therefore lacks an absolute evaluation. Further investigation is needed to examine the correlation between these metric values and human preferences for each generated category. This limitation is part of a broader challenge known as alignment evaluation for LLMs [21]. Finally, we treat the final human-defined categories from the original studies as references, although these may be inconsistent or incomplete.
Internal Validity. The outputs of AstaGen depend on the behavior of the underlying LLM (DeepSeek-V3), which is stochastic by nature. To mitigate variability, we repeated each run ten times. Still, some outcomes may be sensitive to prompt design or batch composition, particularly in Phase 2, where category merging occurs.
External Validity. Our evaluation used SATD datasets from three domains—quantum software, smart contracts, and machine learning software. While these cover diverse domains, they may not generalize to all types of software projects.
@@ VIII Conclusion
We present AstaGen, an initial step toward automating SATD taxonomy generation using LLMs. Across three datasets, AstaGen (i) recovered domain-specific categories reported in prior studies (e.g., Layer Configuration in the ML dataset), (ii) outperformed a naive baseline in comment assignment by up to 0.13 F1, and (iii) completed the entire process for 448 comments in under two hours at a cost of less than one USD even for the largest dataset.
Our findings suggest that fully automating SATD taxonomy generation remains a challenging task. Nevertheless, employing AstaGen as a virtual annotator or as an initial step to propose candidate categories may help reduce human effort. In future work, we plan to explore semi-automated, human-in-the-loop workflows and examine the trade-offs between the degree of automation and taxonomy quality. Furthermore, extending automated taxonomy generation beyond SATD to other open coding tasks (e.g., a taxonomy of root causes of fuzzing build failure [22]) remains an promising direction for future research.
@@ Acknowledgment
We gratefully acknowledge the financial support of: (1) JSPS for the KAKENHI grants (JP24K02921, JP25K03100); (2) Japan Science and Technology Agency (JST) as part of Adopting Sustainable Partnerships for Innovative Research Ecosystem (ASPIRE), Grant NumberJPMJAP2415, and (3) the Inamori Research Institute for Science for supporting Yasutaka Kamei via the InaRIS Fellowship.

### v2

@@ IV-E3 RQ3
We measure the average runtime and DeepSeek-V3 API cost over ten runs per dataset. Each run logs total input/output tokens, and cost is computed using the official DeepSeek-V3 pricing table. [^7] https://api-docs.deepseek.com/quick_start/pricing, as of May 20, 2025
@@ V-A RQ1: Category Name Alignment
Figure 4 shows the distribution of $top\_sim(H_{i})$ scores. The x-axis denotes the dataset, and colors indicate the taxonomy level. The mean top_simtop\_sim scores for the QS, SC, and ML datasets were 0.36, 0.32, and 0.42, respectively.
LLMs can generate domain-specific category names that are similar to human-defined ones. As highlighted in Figure 4, all points with scores $\geq 0.8$ corresponded to domain-specific categories (e.g., Layer Configuration in ML).
High-similarity categories are more likely to be domain-specific, whereas low-similarity categories tend to be generic. To examine the relationship between $top\_sim(H_{i})$ and the type of category, two authors independently labeled each human-defined category in the top and bottom 10% of $top\_sim(H_{i})$ as either domain-specific or generic, counting only those with agreement. Of the top 10% (77 categories), 46 (60%) were domain-specific; in the bottom 10%, 53 (69%) were generic. These results suggest that LLMs are more effective at generating consistent names for domain-specific categories, while generic categories are harder to match semantically.
Answer to RQ1: The pipeline can generate domain-specific category names that are semantically similar to human-defined names. In contrast, low-similarity cases were predominantly generic (69% of 77 categories).
V-B RQ2: Taxonomy Content Alignment
TABLE I: Comparison of the pipeline and its naive variant.
@@ V-B RQ2: Taxonomy Content Alignment
P: The pipeline, N: Naive. Bold indicates the better value.
Table I presents the mean precision, recall, F1 scores, and category granularity for each dataset and taxonomy level. For comparison, we introduced a baseline, the naive variant, which generates the taxonomy directly using an LLM without Phase 1 and 2. This baseline allows us to evaluate the extent to which the pipeline improves alignment with the human-defined taxonomy compared to a naive use of LLMs.
The pipeline achieved higher F1 scores than the naive variant in five out of six cases, indicating a better balance between precision and recall. This result suggests that core components, which the naive variant lacks, are effective in improving the alignment of comment assignments.
It also matched the granularity of human-defined taxonomies more closely in four of six cases, exhibiting a smaller absolute category-count difference than the naive variant. In contrast, the naive variant often produced fewer categories, which likely contributed to its higher recall.
Answer to RQ2: The pipeline outperformed the baseline in F1 (5/6 cases) and generated better category counts (4/6 cases). Precision and recall still lay between 0.152 and 0.638, showing room for improvement.
TABLE II: Average execution time and cost for each dataset.
@@ V-C RQ3: Efficiency (Time and Cost)
Table II shows the average execution time and cost for each dataset, averaged over ten runs.
Compared to prior manual efforts, LLMs achieved significantly reduces both time and cost. For example, Ebrahimi et al. [6] reported spending 57 person-hours just to construct the initial taxonomy, while Xiao et al. [8] required 120 person-hours to manually inspect 500 SATD comments. These substantial efficiency gains highlight the practicality of using LLMs for large scale SATD taxonomy generation.
Answer to RQ3: Laverage LLMs in SATD taxonomy construction is efficient in terms of both time and cost. Even for the largest dataset (i.e., 448 comments), it took only 109.50 minutes and cost just \$0.813 on average.
VI Use Case
While our results indicate that LLMs cannot perfectly replicate human-defined SATD taxonomies, they can meaningfully reduce manual effort. We propose two use cases: (1) as a virtual annotator that can replace some human annotators in collaborative taxonomy construction (e.g., standardizing coding guides) [23] and (2) as an initial step that provides category candidates to support annotators. Notably, LLMs demonstrated the ability to generate domain-specific categories. This makes it especially valuable when researchers must construct a taxonomy in an unfamiliar domain.
@@ VII Threats to Validity
Construct Validity. Our RQ1 and RQ2 evaluations measured only the alignment with the human-defined taxonomy. In RQ1, some LLM-generated categories that did not align could represent novel category discoveries. Future work will investigate whether such categories reflect LLM errors or discoveries of new categories. We also used human-defined taxonomies as references, acknowledging they might be inconsistent or incomplete.
Internal Validity. Results depended on the stochastic behavior of DeepSeek-V3; we averaged ten runs to reduce variance. Outcomes may still vary with prompt design and batch size.
External Validity. Our evaluation used SATD datasets from three domains. While these covered diverse domains, they may not generalize to all types of software projects.
@@ VIII Conclusion
We investigated the extent to which LLMs could generate SATD taxonomies as an initial step toward automated open coding. Across three datasets, our LLM‑driven pipeline (i) recovered domain‑specific categories reported in prior studies (e.g., Layer Configuration in ML), (ii) outperformed the naive baseline by as much as +0.13 F1 in comment assignment, and (iii) completed taxonomy construction for 448 comments in under two hours at a cost below \$1. Although full automation remains challenging, LLMs can reduce human effort. Future work will explore human‑in‑the‑loop pipelines and extend this approach to other open coding tasks beyond SATD taxonomy construction in software engineering.
@@ Acknowledgment
We gratefully acknowledge the financial support of: (1) JSPS for the KAKENHI grants (JP24K02921, JP25K03100, JP25K22845); (2) Japan Science and Technology Agency (JST) as part of Adopting Sustainable Partnerships for Innovative Research Ecosystem (ASPIRE), Grant Number JPMJAP2415, and (3) the Inamori Research Institute for Science for supporting Yasutaka Kamei via the InaRIS Fellowship.